The rules for an essay contest say that entries can have 500 words with an absolute deviation of at most 30 words. Write and solve an absolute value inequality that represents the acceptable numbers of words.
The inequality is ?, where w represents the number of words

Answer by Theo(13342)

the absolute value equation is:

|w-500| <= 30

that says that the absolute value of (w - 500) <= 30

you solve that equation in two ways.

when (w - 500) is positive, the equation becomes w - 500 <= 30.

add 500 to both sides of this equation to get w <= 30 + 500.

simplify to get w <= 530.

when (w - 500) is negative, the equation becomes w - 500 >= -30.

add 500 to both sides of this equation to get w >= -30 + 500.

simplify to get w >= 470.

you have w <= 530 and w >= 470.

this can be written as 470 <= w <= 530.

the maximum deviation is 30 either way.